# MMLongBench-Doc: Benchmarking Long-context Document Understanding with Visualizations

### v1

@@ Abstract
Understanding documents with rich layouts and multi-modal components is a long-standing and practical task. Recent Large Vision-Language Models (LVLMs) have made remarkable strides in various tasks, particularly in single-page document understanding (DU). However, their abilities on long-context DU remain an open problem. This work presents MMLongBench-Docfootnotetext: Project Page: https://mayubo2333.github.io/MMLongBench-Doc, a long-context, multi-modal benchmark comprising 1,062 expert-annotated questions. Distinct from previous datasets, it is constructed upon 130 lengthy PDF-formatted documents with an average of 49.4 pages and 20,971 textual tokens. Towards comprehensive evaluation, answers to these questions rely on pieces of evidence from (1) different sources (text, image, chart, table, and layout structure) and (2) various locations (i.e., page number). Moreover, 33.2% of the questions are cross-page questions requiring evidence across multiple pages. 22.8% of the questions are designed to be unanswerable for detecting potential hallucinations. Experiments on 14 LVLMs demonstrate that long-context DU greatly challenges current models. Notably, the best-performing model, GPT-4o, achieves an F1 score of only 42.7%, while the second-best, GPT-4V, scores 31.4%. Furthermore, 12 LVLMs (all except GPT-4o and GPT-4V) even present worse performance than their LLM counterparts which are fed with lossy-parsed OCR documents. These results validate the necessity of future research toward more capable long-context LVLMs.
@@ 1 Introduction
In this paper, we present MMLongBench-Doc, a benchmark designed to evaluate the Multi-Modality Long-context Document understanding abilities of LVLMs. Towards a comprehensive benchmark, it incorporates lengthy documents from both four existing datasets [13; 17; 18; 19] and other various papers, brochures, etc. Consequently, our benchmark includes 130 PDF-formatted documents spanning across 7 diverse domains, with each document averaging 49.4 pages and 20,970.9 textual tokens. Regarding the questions, we employ ten expert-level annotators to (1) edit questions associated with documents from existing datasets to meet our benchmark’s standard and (2) create new questions for all collected documents to expand the scale of the benchmark. Then a three-round, semi-automatic reviewing process ensures the benchmark’s annotation quality. As a result, MMLongBench-Doc comprises 1,062 human-annotated questions, with 184 sourced from four existing datasets and 878 newly annotated. Being a multi-modal benchmark, the answer to each question requires evidence from one or more of these five in-document sources: text, layout, chart, table, and image. Questions are categorized into three types based on the number of evidence pages  [^1] Given a document $D$ and a question $q$ upon $D$, We call page $P$ (in document $D$) an evidence page of $q$ if the answer of $q$ necessitates one or more pieces of evidence in page $P$., with examples illustrated in Figure 1(a): (1) 467 single-page questions (with one evidence page) mainly to evaluate localization abilities, (2) 353 cross-page questions (with multiple evidence pages) to assess cross-page comprehension, and (3) 242 unanswerable questions (no evidence for answering it, i.e., no evidence pages) to reduce shortcuts and measure LVLMs’ potential hallucinations. Meta-information including evidence pages, sources, and answer formats, is preserved for fine-grained evaluation and analysis. Detailed descriptions of the annotation pipeline and statistics can be found in Section 3.
@@
We conduct extensive experiments on MMLongBench-Doc to evaluate the long-context DU abilities of 14 LVLMs, including 4 proprietary and 10 open-source ones. Given a document, we screenshot each page and feed all of these PNG-formatted images to LVLMs in an end-to-end approach. For comparison, we also convert the documents to textual format by optical character recognition (OCR) and evaluate another 6 proprietary and 4 open-source 10 LLMs (6 proprietary and 4 open-source ones). The results in Figure 1(c) highlight the challenges that current LVLMs face with long-context DU. The best-performing LVLM, GPT-4o, achieves an overall F1 score of only 42.7%, while the second-best LVLM, GPT-4V, scores 31.4%. Moreover, all the remaining LVLMs tested with multi-modal documents performed worse than single-modal LLMs handling lossy, OCR-parsed texts. Specifically, the Gemini-1.5-Pro and Claude-3-Opus present 4.1% and 6.5% absolute decrease when the inputs change from document screenshots to OCR-parsed texts. Regarding open-source models, the best-performing LVLM lags behind the best-performing LLM by 11.9%. These results reveal that long-context DU is a far-from-resolved task for current LVLMs.
@@ 3.1 Document Collection
In summary, we collect a total of 130 documents. Among them, 76 documents are from existing datasets (colored in purple and exemplified as Document 2 in Figure 2, the same below) and incorporate previously annotated questions (represented as triangles). The remaining 54 documents are newly collected (colored in orange and exemplified as Document 1) and incorporate no existing questions. We manually categorize them into 7 types: Research Report, Financial Report, Academic Paper, Brochure, Guideline, Administration & Industry File, Tutorial / Workshop. We showcase some instances of these documents in Appendix A.3.
@@ 3.2 Question and Answer Collection
We employ ten expert-level annotators, including some authors, for the question-and-answer collection. All of them are doctors or Ph.D. students proficient in English reading and writing. Before formal annotation, they undergo a training session and pre-annotate three documents for practice. We iteratively review their annotation results and provide personalized feedback until their annotations meet the standards mentioned above. Regarding the formal annotation, we divide 130 documents into 54 batches (each having 2-4 documents) and dispatch these batches to annotators. We then ask the annotators to submit their results in units of batches and set reasonable time intervals for each batch’s submission. We timely evaluate their annotations after each submission and remind the annotators if their questions in this turn diverge from the standards. It avoids the annotators rushing all assignments in a short time and benefits the annotation quality. We recommend the annotators take 60-90 minutes on each document. Specifically, the annotators shall rapidly read through the whole document in the first 15-30 minutes. For the remaining time, they shall dive deep into specific components to modify existing annotations and/or add new annotations as detailed below.
@@
Add New Questions. We newly annotate questions on both existing and newly collected documents to expand the questions in our benchmark. See square-marked questions in Figure 2 as examples. Specifically, we ask annotators to add about 3 questions on existing documents, and 6 questions on newly-collected documents. Given most existing questions (even after editing) are single-page ones and sourced from texts, we put more focus on (1) cross-page and unanswerable questions and (2) questions sourced from tables, charts, and images for newly added questions to balance the distribution. We detail the quantitative requirements in Appendix A.5. Associated with questions, annotators also provide reference answers and meta-information (i.e., evidence sources, answer format, evidence locations) for all samples. We finalized a collection of 936 samples in this procedure. The corresponding GUI is shown in Appendix A.6.
@@ 3.3 Quality Control
- Average/Medium pages 49.4 / 31
- Average/Medium length 20,970.9 / 12,146
Total questions 1,062
- Single-page question 467 (44.0%)
- Cross-page questions 353 (33.2%)
- Unanswerable questions 242 (22.8%)
- Derived questions 184 (17.3%)
- Newly-annotated questions 878 (82.7%)
@@
- Pure-text 296 (36.1%)
- Layout 119 (14.5%)
- Table 212 (25.9%)
- Chart 168 (20.5%)
- Image 282 (34.4%)
@@
- String 239 (29.2%)
- Integer 280 (34.1%)
- Float 158 (19.3%)
- List 143 (17.4%)
Avg./Max. question length 16.6 / 60
@@ 3.4 Dataset Overview and Analysis
The main statistics of MMLongBench-Doc are presented in Table 2. Overall, our benchmark consists of 1,062 questions. These questions are constructed upon 130 lengthy documents across 7 document types, with an average of 49.4 pages and 20,970.9 tokens. Please see detailed distributions of these documents in Figure 3. Regarding the questions, there are 415 single-page questions (1 evidence page), 353 cross-page questions (2+ evidence pages), and 242 unanswerable questions (no evidence page). These three types of questions evaluate the LVLMs’s long-context DU capabilities from complementary aspects: the localization ability, the cross-page comprehension ability, and the hallucination severity, respectively. For single-page and cross-page questions, their answer evidence is scattered among different context sources (i.e., text, layout, table, chart, image) and evenly distributed across different locations of the documents (see Table 2, Figure 4 Left and Middle). Also notably, 29.2% of cross-page questions have more than two evidence pages, which further enhances the challenge of our benchmark.
@@ 4.2 Experimental Setup
We evaluate 14 LVLMs on MMLongBench-Doc, including 4 proprietary LVLMs and 10 open-source LVLMs. To purely evaluate LVLMs’ long-context DU abilities, we screenshot each page of the PDF-formatted document with 120 DPI and feed all these PNG-formatted images to LVLMs in an end-to-end approach. Notably, all evaluated open-source LVLMs do not support multi-image inputs or present significant performance drops when fed with excessive images (e.g., more than 10 or 20 images). Therefore, we employ a concatenation strategy that combines all screenshot pages into 1 or 5 images and feeds these concatenated images to open-source LVLMs. Regarding proprietary LVLMs, we adopt the same concatenation strategy and reduce the image number to 20 for Claude-3-Opus to fit its maximum image threshold. For GPT-4o, GPT-4V, and Gemini-1.5-Pro, we directly send all original screenshots to them (i.e., the image number equals the page number).
@@
ChatGLM-128k [37] 6B 128k 25.0 12.7 9.1 14.3 12.1 20.4 11.9 18.2 17.1 15.7
Mistral-Instruct-v0.2 [44] 7B 32k 21.2 13.4 10.2 11.1 10.6 17.1 11.4 25.2 16.9 13.9
Mixtral-Instruct-v0.1 [45] 8x7B 32k 26.2 15.6 11.9 17.4 13.3 22.7 13.9 13.6 17.7 17.6
Mixtral-Instruct-v0.1 [45] 8x22B 64k 34.8 20.5 19.6 24.2 18.5 28.4 22.4 32.6 27.3 25.0
@@
QWen-Plus [46] - 32k 18.1 16.4 6.9 8.4 8.6 14.2 10.8 41.7 19.2 13.6
DeepSeek-V2 [47] - 32k 28.0 20.5 9.4 17.6 9.7 20.9 15.4 47.5 25.1 19.8
Claude-3 Opus [4] - 32k 30.9 32.3 14.4 28.4 17.5 33.8 18.4 30.3 27.4 25.1
Gemini-1.5-Pro [3] - 32k 29.5 15.1 13.1 18.6 10.8 22.5 15.6 74.0 31.8 25.2
GPT-4-turbo [48] - 128k 37.8 21.0 21.5 29.1 16.8 31.5 23.4 31.4 28.6 27.0
GPT-4o [2] - 128k 43.7 23.3 28.1 38.1 22.4 37.9 30.6 18.2 31.1 31.9
@@
Claude-3 Opus [4] - 200k 27.7 24.7 14.8 13.2 17.7 27.4 13.5 7.4 16.7 18.6
Gemini-1.5-Pro [3] - 128k 20.6 17.9 5.7 16.4 15.2 21.4 11.3 69.4 28.8 20.9
GPT-4V(ision) [48] - 128k 34.8 28.3 28.4 35.2 26.7 37.4 27.2 30.6 32.5 31.4
GPT-4o [2] - 128k 43.6 43.5 43.1 49.9 41.6 54.0 37.5 19.8 40.8 42.7
@@ 4.3 Main Results
We compare the performance of different LVLMs and LLMs in Table 3, reporting their generalized accuracy and F1 scores (shown in the last two columns). Regarding LVLMs, we draw several conclusions as below: (1) The performance demonstrates that long-context DU is still a challenging and unsolved task for current LVLMs. The best-performing LVLM, GPT-4o, merely achieves a 42.7% F1 score. The second best-performing LVLM, GPT-4V, lags behind by over 10% percent and presents a 31.4% F1 score. All other LVLMs only achieve about 20% or even lower F1 scores. (2) Though far from satisfactory, GPT-4o performs much better than all other models (including GPT-4V). Thus we speculate that the multi-modal pre-training paradigm significantly benefits LVLMs’ cross-modality understanding capabilities. (3) Proprietary LVLMs perform better than open-source LVLMs by a large margin. We attribute it to the difference of acceptable image numbers: open-source LVLMs only support single-image or several-image  [^6] We observe that all open-source LVLMs struggle when fed with more than 5 images. inputs, while proprietary LVLMs can be fed with at least 20 images or even more. Given that lengthy documents have tens of even hundreds of pages, it is impractical for open-source LVLMs to accurately perceive the information in the documents from the excessively concatenated images. (4) The performances of different models are highly correlated with their acceptable image numbers and maximum image resolutions. Notably, open-source LVLMs that support high-resolution images (i.e., InternLM-XC2-4KHD and InternVL-Chat-v1.5) exhibit superior performance compared to those with lower resolution limits.
Surprisingly, LVLMs even demonstrate overall worse performance than LLMs, even LLMs are fed with lossy OCR-parsed documents. Specifically, Gemini-1.5-Pro and Claude-3 Opus have 4.3% and 6.5% absolute F1-score degradations on vision versions. And the best-performing LLM (Mixtral) also surpasses the best-performing LVLM (InternVL-v1.5) by 11.9%. The above results clearly reveal that most current LVLMs are still not proficient in cross-modality, long-context document understandings. It is promising that GPT-4o and GPT-4-turbo achieve better performance when seeing multi-modality PDF documents than parsed text by 10.8% and 4.4% F1-score, respectively. Their performances validate the feasibility, benefit, and necessity of understanding documents in an end-to-end, cross-modality approach. We speculate that the scarce related pre-training corpus (i.e., extremely multi-image or lengthy documents) hinders the long-context DU capabilities of other LVLMs. We will leave related explorations for future work.

### v2

@@ Abstract
Understanding documents with rich layouts and multi-modal components is a long-standing and practical task. Recent Large Vision-Language Models (LVLMs) have made remarkable strides in various tasks, particularly in single-page document understanding (DU). However, their abilities on long-context DU remain an open problem. This work presents MMLongBench-Docfootnotetext: Project Page: https://mayubo2333.github.io/MMLongBench-Doc, a long-context, multi-modal benchmark comprising 1,091 expert-annotated questions. Distinct from previous datasets, it is constructed upon 135 lengthy PDF-formatted documents with an average of 47.5 pages and 21,214 textual tokens. Towards comprehensive evaluation, answers to these questions rely on pieces of evidence from (1) different sources (text, image, chart, table, and layout structure) and (2) various locations (i.e., page number). Moreover, 33.0% of the questions are cross-page questions requiring evidence across multiple pages. 22.5% of the questions are designed to be unanswerable for detecting potential hallucinations. Experiments on 14 LVLMs demonstrate that long-context DU greatly challenges current models. Notably, the best-performing model, GPT-4o, achieves an F1 score of only 44.9%, while the second-best, GPT-4V, scores 30.5%. Furthermore, 12 LVLMs (all except GPT-4o and GPT-4V) even present worse performance than their LLM counterparts which are fed with lossy-parsed OCR documents. These results validate the necessity of future research toward more capable long-context LVLMs.
@@ 1 Introduction
In this paper, we present MMLongBench-Doc, a benchmark designed to evaluate the Multi-Modality Long-context Document understanding abilities of LVLMs. Towards a comprehensive benchmark, it incorporates lengthy documents from both four existing datasets [13; 17; 18; 19] and other various papers, brochures, etc. Consequently, our benchmark includes 135 PDF-formatted documents spanning across 7 diverse domains, with each document averaging 47.5 pages and 21,214.1 textual tokens. Regarding the questions, we employ ten expert-level annotators to (1) edit questions associated with documents from existing datasets to meet our benchmark’s standard and (2) create new questions for all collected documents to expand the scale of the benchmark. Then a three-round, semi-automatic reviewing process ensures the benchmark’s annotation quality. As a result, MMLongBench-Doc comprises 1,091 human-annotated questions, with 184 sourced from four existing datasets and 907 newly annotated. Being a multi-modal benchmark, the answer to each question requires evidence from one or more of these five in-document sources: text, layout, chart, table, and image. Questions are categorized into three types based on the number of evidence pages  [^1] Given a document $D$ and a question $q$ upon $D$, We call page $P$ (in document $D$) an evidence page of $q$ if the answer of $q$ necessitates one or more pieces of evidence in page $P$., with examples illustrated in Figure 1(a): (1) 485 single-page questions (with one evidence page) mainly to evaluate localization abilities, (2) 360 cross-page questions (with multiple evidence pages) to assess cross-page comprehension, and (3) 246 unanswerable questions (no evidence for answering it, i.e., no evidence pages) to reduce shortcuts and measure LVLMs’ potential hallucinations. Meta-information including evidence pages, sources, and answer formats, is preserved for fine-grained evaluation and analysis. Detailed descriptions of the annotation pipeline and statistics can be found in Section 3.
@@
We conduct extensive experiments on MMLongBench-Doc to evaluate the long-context DU abilities of 14 LVLMs, including 4 proprietary and 10 open-source ones. Given a document, we screenshot each page and feed all of these PNG-formatted images to LVLMs in an end-to-end approach. For comparison, we also convert the documents to textual format by optical character recognition (OCR) and evaluate another 6 proprietary and 4 open-source 10 LLMs (6 proprietary and 4 open-source ones). The results in Figure 1(c) highlight the challenges that current LVLMs face with long-context DU. The best-performing LVLM, GPT-4o, achieves an overall F1 score of only 44.9%, while the second-best LVLM, GPT-4V, scores 30.5%. Moreover, all the remaining LVLMs tested with multi-modal documents performed worse than single-modal LLMs handling lossy, OCR-parsed texts. Specifically, the Gemini-1.5-Pro and Claude-3-Opus present 4.2% and 6.4% absolute decrease when the inputs change from document screenshots to OCR-parsed texts. Regarding open-source models, the best-performing LVLM lags behind the best-performing LLM by 11.7%. These results reveal that long-context DU is a far-from-resolved task for current LVLMs.
@@ 3.1 Document Collection
In summary, we collect a total of 135 documents. Among them, 76 documents are from existing datasets (colored in purple and exemplified as Document 2 in Figure 2, the same below) and incorporate previously annotated questions (represented as triangles). The remaining 59 documents are newly collected (colored in orange and exemplified as Document 1) and incorporate no existing questions. We manually categorize them into 7 types: Research Report, Financial Report, Academic Paper, Brochure, Guideline, Administration & Industry File, Tutorial / Workshop. We showcase some instances of these documents in Appendix A.3.
@@ 3.2 Question and Answer Collection
We employ ten expert-level annotators, including some authors, for the question-and-answer collection. All of them are doctors or Ph.D. students proficient in English reading and writing. Before formal annotation, they undergo a training session and pre-annotate three documents for practice. We iteratively review their annotation results and provide personalized feedback until their annotations meet the standards mentioned above. Regarding the formal annotation, we divide 135 documents into 54 batches (each having 2-4 documents) and dispatch these batches to annotators. We then ask the annotators to submit their results in units of batches and set reasonable time intervals for each batch’s submission. We timely evaluate their annotations after each submission and remind the annotators if their questions in this turn diverge from the standards. It avoids the annotators rushing all assignments in a short time and benefits the annotation quality. We recommend the annotators take 60-90 minutes on each document. Specifically, the annotators shall rapidly read through the whole document in the first 15-30 minutes. For the remaining time, they shall dive deep into specific components to modify existing annotations and/or add new annotations as detailed below.
@@
Add New Questions. We newly annotate questions on both existing and newly collected documents to expand the questions in our benchmark. See square-marked questions in Figure 2 as examples. Specifically, we ask annotators to add about 3 questions on existing documents, and 6 questions on newly-collected documents. Given most existing questions (even after editing) are single-page ones and sourced from texts, we put more focus on (1) cross-page and unanswerable questions and (2) questions sourced from tables, charts, and images for newly added questions to balance the distribution. We detail the quantitative requirements in Appendix A.5. Associated with questions, annotators also provide reference answers and meta-information (i.e., evidence sources, answer format, evidence locations) for all samples. We finalized a collection of 965 samples in this procedure. The corresponding GUI is shown in Appendix A.6.
@@ 3.3 Quality Control
- Average/Medium pages 47.5 / 28
- Average/Medium length 21,214.1 / 12,179
Total questions 1,091
- Single-page question 485 (44.5%)
- Cross-page questions 360 (33.0%)
- Unanswerable questions 246 (22.5%)
- Derived questions 184 (16.9%)
- Newly-annotated questions 907 (83.1%)
@@
- Pure-text 298 (35.3%)
- Layout 119 (14.1%)
- Table 218 (25.8%)
- Chart 178 (21.1%)
- Image 293 (34.7%)
@@
- String 250 (29.5%)
- Integer 290 (34.2%)
- Float 160 (18.9%)
- List 147 (17.3%)
Avg./Max. question length 16.4 / 60
@@ 3.4 Dataset Overview and Analysis
The main statistics of MMLongBench-Doc are presented in Table 2. Overall, our benchmark consists of 1,091 questions. These questions are constructed upon 135 lengthy documents across 7 document types, with an average of 47.5 pages and 21,214.1 tokens. Please see detailed distributions of these documents in Figure 3. Regarding the questions, there are 485 single-page questions (1 evidence page), 360 cross-page questions (2+ evidence pages), and 246 unanswerable questions (no evidence page). These three types of questions evaluate the LVLMs’s long-context DU capabilities from complementary aspects: the localization ability, the cross-page comprehension ability, and the hallucination severity, respectively. For single-page and cross-page questions, their answer evidence is scattered among different context sources (i.e., text, layout, table, chart, image) and evenly distributed across different locations of the documents (see Table 2, Figure 4 Left and Middle). Also notably, 29.0% of cross-page questions have more than two evidence pages, which further enhances the challenge of our benchmark.
@@ 4.2 Experimental Setup
We evaluate 14 LVLMs on MMLongBench-Doc, including 4 proprietary LVLMs and 10 open-source LVLMs. To purely evaluate LVLMs’ long-context DU abilities, we screenshot each page of the PDF-formatted document with 144 DPI and feed all these PNG-formatted images to LVLMs in an end-to-end approach. Notably, all evaluated open-source LVLMs do not support multi-image inputs or present significant performance drops when fed with excessive images (e.g., more than 10 or 20 images). Therefore, we employ a concatenation strategy that combines all screenshot pages into 1 or 5 images and feeds these concatenated images to open-source LVLMs. Regarding proprietary LVLMs, we adopt the same concatenation strategy and reduce the image number to 20 for Claude-3-Opus to fit its maximum image threshold. For GPT-4o, GPT-4V, and Gemini-1.5-Pro, we directly send all original screenshots to them (i.e., the image number equals the page number).
@@
ChatGLM-128k [37] 6B 128k 23.4 12.7 9.7 10.2 12.2 18.8 11.5 18.1 16.3 14.9
Mistral-Instruct-v0.2 [44] 7B 32k 19.9 13.4 10.2 10.1 11.0 16.9 11.3 24.1 16.4 13.8
Mixtral-Instruct-v0.1 [45] 8x7B 32k 24.2 14.8 12.5 15.0 13.7 21.3 14.1 13.1 17.0 16.9
Mixtral-Instruct-v0.1 [45] 8x22B 64k 34.2 21.3 19.5 21.3 19.2 27.7 21.9 32.4 26.9 24.7
@@
QWen-Plus [46] - 32k 17.4 15.6 7.4 7.9 8.8 14.2 10.6 42.2 18.9 13.4
DeepSeek-V2 [47] - 32k 27.8 19.6 8.8 17.0 9.4 20.2 15.4 48.1 24.9 19.6
Claude-3 Opus [4] - 32k 30.8 30.1 16.4 24.4 16.3 32.0 18.6 30.9 26.9 24.5
Gemini-1.5-Pro [3] - 32k 29.3 15.9 12.5 17.7 11.5 21.2 16.4 73.4 31.2 24.8
GPT-4-turbo [48] - 128k 36.5 21.0 20.7 24.3 17.3 28.7 23.8 31.2 27.6 25.9
GPT-4o [2] - 128k 41.1 23.4 28.5 38.1 22.4 35.4 29.3 18.6 30.1 30.5
@@
Claude-3 Opus [4] - 200k 24.9 24.7 14.8 13.0 17.1 25.6 13.8 7.6 17.4 18.1
Gemini-1.5-Pro [3] - 128k 21.0 17.6 6.9 14.5 15.2 21.1 11.1 69.2 28.2 20.6
GPT-4V(ision) [48] - 128k 34.4 28.3 28.2 32.4 26.8 36.4 27.0 31.2 32.4 31.2
GPT-4o [2] - 128k 46.3 46.0 45.3 50.0 44.1 54.5 41.5 20.2 42.8 44.9
@@ 4.3 Main Results
We compare the performance of different LVLMs and LLMs in Table 3, reporting their generalized accuracy and F1 scores (shown in the last two columns). Regarding LVLMs, we draw several conclusions as below: (1) The performance demonstrates that long-context DU is still a challenging and unsolved task for current LVLMs. The best-performing LVLM, GPT-4o, merely achieves a 44.9% F1 score. The second best-performing LVLM, GPT-4V, lags behind by over 10% percent and presents a 31.4% F1 score. All other LVLMs only achieve about 20% or even lower F1 scores. (2) Though far from satisfactory, GPT-4o performs much better than all other models (including GPT-4V). Thus we speculate that the multi-modal pre-training paradigm significantly benefits LVLMs’ cross-modality understanding capabilities. (3) Proprietary LVLMs perform better than open-source LVLMs by a large margin. We attribute it to the difference of acceptable image numbers: open-source LVLMs only support single-image or several-image  [^6] We observe that all open-source LVLMs struggle when fed with more than 5 images. inputs, while proprietary LVLMs can be fed with at least 20 images or even more. Given that lengthy documents have tens of even hundreds of pages, it is impractical for open-source LVLMs to accurately perceive the information in the documents from the excessively concatenated images. (4) The performances of different models are highly correlated with their acceptable image numbers and maximum image resolutions. Notably, open-source LVLMs that support high-resolution images (i.e., InternLM-XC2-4KHD and InternVL-Chat-v1.5) exhibit superior performance compared to those with lower resolution limits.
Surprisingly, LVLMs even demonstrate overall worse performance than LLMs, even LLMs are fed with lossy OCR-parsed documents. Specifically, Gemini-1.5-Pro and Claude-3 Opus have 4.2% and 6.4% absolute F1-score degradations on vision versions. And the best-performing LLM (Mixtral) also surpasses the best-performing LVLM (InternVL-v1.5) by 11.7%. The above results clearly reveal that most current LVLMs are still not proficient in cross-modality, long-context document understandings. It is promising that GPT-4o and GPT-4-turbo achieve better performance when seeing multi-modality PDF documents than parsed text by 14.4% and 5.3% F1-score, respectively. Their performances validate the feasibility, benefit, and necessity of understanding documents in an end-to-end, cross-modality approach. We speculate that the scarce related pre-training corpus (i.e., extremely multi-image or lengthy documents) hinders the long-context DU capabilities of other LVLMs. We will leave related explorations for future work.